L I S T

of the scientific publications of

Bozhidar Zakhariev ILIEV^{∗}^{†} ^{‡}

⇓ \underline{1981} ⟹ • 1981 · ⟶ {↙ \overleftarrow{1981} \atop ↓}

1. Iliev B.Z. and S. Manov, Deviation equations in spaces with torsion, → In: Proceedings of the 5-th Soviet (USSR) Gravity Conference “Modern theoretical and experimental problems of relativity theory and gravitation”, Moscow Univ., Moscow, 1981, p. 122, In Russian. → http://arXiv.org e-Print archive, E-print No. gr-qc/0507002, July 1, 2005, In English.

⇓ \underline{1983} ⟹ • 1983 · ⟶ {↙ \overleftarrow{1983} \atop ↓}

2. Iliev B.Z. and S. Manov, Deviation equations in spaces with affine connection, → Communication JINR, P2-83-897, Dubna, 1983, 14 pp., In Russian → http://arXiv.org e-Print archive, E-print No. math.DG/0512008, December 1, 2005, In English.

⇓ \underline{1984} ⟹ • 1984 · ⟶ {↙ \overleftarrow{1984} \atop ↓}

3. Iliev B.Z., Deviation equations and the weak equivalence principle in spaces with affine connection, → In: Proceedings of the 6-th Soviet (USSR) Gravity Conference “Modern theoretical and experimental problems of relativity theory and gravitation”, University ”Friendship between nations”, Moscow 3-5 July 1984, University ”Friendship between nations” Press, Moscow, 1984, pp. 227, In Russian. → http://arXiv.org e-Print archive, E-print No. gr-qc/0605008, May 1, 2006, In English.

4. Iliev B.Z., The deviation equation and the equivalence principle in spaces with affine connection, → Comptes rendus de L’Academie bulgare des Sciences, vol. 37, No. 9, 1984, pp. 1187–1190.

⇓ \underline{1986} ⟹ • 1986 · ⟶ {↙ \overleftarrow{1986} \atop ↓}

5. Iliev B.Z., Deviation equations in spaces with affine connection. I. Generalized deviation equation: physical interpretation, local case and nonlocal problems. II. Displacement vector: the general case, → Bulgarian Journal of Physics, vol. 13, No. 6, 1986, pp. 494–506.

∗Laboratory of Mathematical Modeling in Physics, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, blvd. Tzarigradsko Chaussée 72, 1784 Sofia, Bulgaria, tel.:+359-2-979-5646, fax:+359-2-9753619.

†E-mail address: bozho@inrne.bas.bg

‡URL: http://theo.inrne.bas.bg/∼bozho/

\underline{List of the scientific publications of Bozhidar Z. Iliev} 2

⇓ \underline{1987} ⟹ • 1987 · ⟶ {↙ \overleftarrow{1987} \atop ↓}

6. Iliev B.Z., Deviation equations in spaces with affine connection. III. Displacement vector: the case of a space with affine connection. IV. Nonlocal deviation equation, → Bulgarian Journal of Physics, vol. 14, No. 1, 1987, pp. 18–31.

7. Iliev B.Z., The deviation equation as an equation of motion, → Communication JINR, E2-87-78, Dubna, 1987, 11 pp.

8. Iliev B.Z., Deviation equations in spaces with affine connection. V. Some approximate methods, → Bulgarian Journal of Physics, vol. 14, No. 3, 1987, pp. 215–220.

9. Iliev B.Z., Linear transports of tensors along curves: General S-transport, → Comptes rendus de L’Academie bulgare des Sciences, vol. 40, No. 7, 1987, pp. 47–50.

10. Iliev B.Z., General linear transport (I-transport), → Comptes rendus de L’Academie bulgare des Sciences, vol. 40, No. 8, 1987, pp. 45–48.

11. Iliev B.Z., Generalized transports and displacement vectors, → Communication JINR, E2-87-267, Dubna, 1987, 9 pp.

⇓ \underline{1989} ⟹ • 1989 · ⟶ {↙ \overleftarrow{1989} \atop ↓}

12. Iliev B.Z., Consistency between metrics and linear transports along curves, → Comptes rendus de L’Academie bulgare des Sciences, vol. 42, No. 3, 1989, pp. 17–20.

13. Iliev B.Z., Relative velocity, momentum and energy of point particles in spaces with general linear transport, → Communication JINR, E2-89-616, Dubna, 1989, 16 pp.

⇓ \underline{1990} ⟹ • 1990 · ⟶ {↙ \overleftarrow{1990} \atop ↓}

14. Iliev B.Z., On the first order deviation equations and some quantities connected with them, → Communication JINR, E2-90-534, Dubna, 1990, 17 pp.

⇓ \underline{1991} ⟹ • 1991 · ⟶ {↙ \overleftarrow{1991} \atop ↓}

15. Iliev B.Z., On one generalization of the Doppler effect in spaces with general linear transport, → In: Proceedings of the 3-d seminar “Gravitational energy and gravitational waves”, JINR, Dubna, 19–21 may 1990, JINR P2-91-164, Dubna, 1991, pp. 96–104, In Russian.

⇓ \underline{1992} ⟹ • 1992 · ⟶ {↙ \overleftarrow{1992} \atop ↓}

16. Iliev B.Z., On some generalizations of the Jacobi identity. I., → Bulletin de la Société des Sciences et des Lettres de Łodz, vol. 42, No. 21, Série: Recherches sur les Déformations, vol. 14, No. 131, pp. 5–11, 1992.

17. Iliev B.Z., On some generalizations of the Jacobi identity. II., → Bulletin de la Société des Sciences et des Lettres de Łodz, vol. 42, No. 28, Série: Recherches sur les Déformations, vol. 14, No. 138, pp. 93-102, 1992; → http://arXiv.org e-Print archive, E-print No. math.DG/0306019, June 18. Iliev B.Z., Consistency between metrics and linear transports along curves,

→ Communication JINR, E5-92-486, Dubna, 1992, 23 pp.

19. Iliev B.Z., Special bases for derivations of tensor algebras. I. Cases in a neighborhood and at a point, → Communication JINR, E5-92-507, Dubna, 1992, 19 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0303373, March 2003.

20. Iliev B.Z., Special bases for derivations of tensor algebras. II. Case along paths, → Communication JINR, E5-92-508, Dubna, 1992, 16 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0304157, April 2003.

21. Iliev B.Z., Special bases for derivations of tensor algebras. III. Case along smooth maps with separable points of selfintersection, → Communication JINR, E5-92-543, Dubna, 1992, 15 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0305061, May 2003.

22. Iliev B.Z., Flat linear connections in terms of flat linear transports in tensor bundles, → Communication JINR, E5-92-544, Dubna, 1992, 20 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0406010, June 1, 2004.

⇓ \underline{1993} ⟹ • 1993 · ⟶ {↙ \overleftarrow{1993} \atop ↓}

23. Iliev B.Z., The equivalence principle in spaces with a linear transport along paths., → In: Proceedings of the 5-th seminar “Gravitational energy and gravitational waves”, JINR, Dubna, 16–18 may 1992, JINR P2-92-559, Dubna, 1993, pp. 69–84, In Russian.

24. Iliev B.Z., Parallel transports in tensor spaces generated by derivations of tensor algebras, → Communication JINR, E5-93-1, Dubna, 1993, 19 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0502009, February 1, 2005.

25. Iliev B.Z., Linear transports along paths in vector bundles. I. General theory, → Communication JINR, E5-93-239, Dubna, 1993, 19 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0411023, November 1, 2004.

26. Iliev B.Z., Linear transports along paths in vector bundles. II. Some applications, → Communication JINR, E5-93-260, Dubna, 1993, 24 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0412010, December 1, 2004.

27. Iliev B.Z., Linear transports along paths in vector bundles. III. Curvature and torsion, → Communication JINR, E5-93-261, Dubna, 1993, 12 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0502008, February 1, 2005.

28. Iliev B.Z., Transports along paths in fibre bundles. General theory, → Communication JINR, E5-93-299, Dubna, 1993, 26 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0503005, March 1, 2005.

⇓ \underline{1994} ⟹ • 1994 · ⟶ {↙ \overleftarrow{1994} \atop ↓}

29. Iliev B.Z., Some generalizations of the Jacobi identity with applications to the curvature- and torsion-depending hamiltonians of particle systems, → In: Deformations of Mathematical Structures II. Hurwitz-type structures and applications to surface physics, Papers from the Seminars on Deformations 1988-1992, Ed. J. Ławrynowicz, Kluwer, Academic Publishers Group, Dordrecht-Boston-London, pp. 161–188, 1994.

30. Iliev B.Z., Linear transports along paths in vector bundles. IV. Consistency with bundle metrics, → Communication JINR, E5-94-17, Dubna, 1994, 22 p; → http://arXiv.org e-Print archive, E-print No. math.DG/0505008, May 1, 2005.

31. Iliev B.Z., Transports along paths in fibre bundles II. Ties with the theory of connections and parallel transports, → Communication JINR, E5-94-16, Dubna, 1994, 24 pp.; → http://arXiv.org e-Print archive, E-print No. math.DG/0503005, March 1, 2005.

32. Iliev B.Z., Transports along paths in fibre bundles III. Consistency with bundle morphisms, → Communication JINR, E5-94-41, Dubna, 1994, 19 pp.; → http://arXiv.org e-Print archive, E-print No. dg-ga/9704004, April 1997.

33. Iliev B.Z., Deviation equations in spaces with a transport along paths, → Communication JINR, E2-94-40, Dubna, 1994, 22 pp.; → http://arXiv.org e-Print archive, E-print No. math-ph/0303002, March 2003.

34. Iliev B.Z., Relative mechanical quantities in spaces with a transport along paths, → Communication JINR, E2-94-188, Dubna, 1994, 19 pp.; → http://arXiv.org e-Print archive, E-print No. math-ph/0310024, October 2003.

35. Iliev B.Z., First order deviation equations in spaces with a transport along paths, → Communication JINR, E2-94-189, Dubna, 1994, 15 pp.; → http://arXiv.org e-Print archive, E-print No. math-ph/0303038, March 2003.

⇓ \underline{1995} ⟹ • 1995 · ⟶ {↙ \overleftarrow{1995} \atop ↓}

36. Iliev B.Z., Centre of mass in spaces with torsion free flat linear connection, → Communication JINR, E5-95-159, Dubna, 1995, 15 pp.; → http://arXiv.org e-Print archive, E-print No. gr-qc/0404002, April 2004.

37. Iliev B.Z., Generalized Doppler effect in spaces with a transport along paths, → Communication JINR, E5-95-160, Dubna, 1995, 12 pp.; → http://arXiv.org e-Print archive, E-print No. math-ph/0401033, January

⇓ \underline{1996} ⟹ • 1996 · ⟶ {↙ \overleftarrow{1996} \atop ↓}

38. Iliev B.Z., Normal frames and the validity of the equivalence principle: I. Cases in a neighborhood and at a point, → Journal of Physics A: Mathematical and General, vol. 29, No. 21, pp. 6895– 6901, 1996. DOI No. 10.1088/0305-4470/29/21/020; → http://arXiv.org e-Print archive, E-print No. gr-qc/9608019, August 1996.

⇓ \underline{1997} ⟹ • 1997 · ⟶ {↙ \overleftarrow{1997} \atop ↓}

39. Iliev B.Z., Linear transports along paths in vector bundles. V. Properties of curvature and torsion, → Communication JINR, E5-97-1, Dubna, 1997, 11 pp.; → http://arXiv.org e-Print archive, E-print No. dg-ga/9709017, September 1997.

40. Iliev B.Z., Transports along maps in fibre bundles, → Communication JINR, E5-97-2, Dubna, 1997, 19 pp.; → http://arXiv.org e-Print archive, E-print No. dg-ga/9709016, September 1997.

41. Iliev B.Z., Normal frames and the validity of the equivalence principle II. The case along paths, → Journal of Physics A: Mathematical and General, vol. 30, No. 12 pp. 4327– 4336, 1997. DOI No.10.1088/0305-4470/30/12/019; → http://arXiv.org e-Print archive, E-print No. gr-qc/9709053, September 1997.

⇓ \underline{1998} ⟹ • 1998 · ⟶ {↙ \overleftarrow{1998} \atop ↓}

42. Iliev B.Z., Is the principle of equivalence a principle?

→ Journal of Geometry and Physics, vol. 24, No. 3, pp. 209–222, 1998. DOI No. 10.1016/S0393-0440(97)00011-9;

https://www.sciencedirect.com/science/article/pii/S0393044097000119; → http://arXiv.org e-Print archive, E-print No. gr-qc/9806062, June 1998.

43. Iliev B.Z., The principle of equivalence: axiom or a theorem?, → In: Proceedings of the 9-th international seminar “Gravitational energy and gravitational waves”, JINR, Dubna, 8–12 December 1996, JINR P2-97-401, Dubna, 1998, pp. 97–104, In Russian.

44. Iliev B.Z., Normal frames and the validity of the equivalence principle III. The case along smooth maps with separable points of self-intersection, → Journal of Physics A: Mathematical and General, vol. 31, No. 4, pp. 1287– 1296, 1998. DOI No. 10.1088/0305-4470/31/4/016; → http://arXiv.org e-Print archive, E-print No. gr-qc/9805088, May 1998.

45. Iliev B.Z., Quantum mechanics from a geometric-observer’s viewpoint, → Journal of Physics A: Mathematical and General, vol. 31, No. 4, pp. 1297– 1305, 1998. DOI No. 10.1088/0305-4470/31/4/017; → http://arXiv.org e-Print archive, E-print No. quant-ph/9803083, March

⇓ \underline{1999} ⟹ • 1999 · ⟶ {↙ \overleftarrow{1999} \atop ↓}

46. Iliev B.Z., Comments on: “Quantum mechanics as a gauge theory of metaplectic spinor fields” by M. Reuter [Int. J. Mod. Phys. vol. A13 (1998), 3835–3884; http://www.arXiv.org e-Print archive, E-print No. hep-th/9804036], → International Journal of Modern Physics A, vol. 14, No. 26, pp. 4153– 4159, 1999. DOI No. 10.1142/S0217751X99001949; → http://arXiv.org e-Print archive, E-print No. quant-ph/9902067, February 1999.

⇓ \underline{2000} ⟹ • 2000 · ⟶ {↙ \overleftarrow{2000} \atop ↓}

47. Iliev B.Z., On a connection between Lorentzian and Euclidean metrics, → Journal of Geometry and Physics, vol. 34, No. 4, pp. 321–335, 2000. DOI No. 10.1016/S0393-0440(99)00074-1; → http://arXiv.org e-Print archive, E-print No. gr-qc/9802057, February 1998.

48. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics (full version), → http://arXiv.org e-Print archive, E-print No. quant-ph/0004041, April 2000.

⇓ \underline{2001} ⟹ • 2001 · ⟶ {↙ \overleftarrow{2001} \atop ↓}

49. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics. I. Introduction. The evolution transport, → Journal of Physics A: Mathematical and General, vol. 34, No. 23, pp. 4887– 4918, 2001. DOI No. 10.1088/0305-4470/34/23/308; → http://arXiv.org e-Print archive, E-print No. quant-ph/9803084, March 1998.

50. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics. II. Equations of motion and observables, → Journal of Physics A: Mathematical and General, vol. 34, No. 23, pp. 4919– 4934, 2001. DOI No. 10.1088/0305-4470/34/23/309; https://iopscience.iop.org/article/10.1088/0305-4470/34/23/309; → http://arXiv.org e-Print archive, E-print No. quant-ph/9804062, April 1998.

51. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics. III. Pictures and integrals of motion, → Journal of Physics A: Mathematical and General, vol. 34, No. 23, pp. 4935– 4950, 2001. DOI No. 10.1088/0305-4470/34/23/310; → http://arXiv.org e-Print archive, E-print No. quant-ph/9806046, June 1998.

⇓ \underline{2002} ⟹ • 2002 · ⟶ {↙ \overleftarrow{2002} \atop ↓}

52. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics. IV. Mixed states and evolution transport’s curvature, → International Journal of Modern Physics A, vol. 17, No. 2, pp. 229–243, 2002. DOI No. 10.1142/S0217751X02005669; → http://arXiv.org e-Print archive, E-print No. quant-ph/9901039, January 53. Iliev B.Z., Fibre bundle formulation of nonrelativistic quantum mechanics. V.

Interpretation, summary, and discussion, → International Journal of Modern Physics A, vol. 17, No. 2, pp. 245–258, 2002. DOI No. 10.1142/S0217751X02005712; → http://arXiv.org e-Print archive, E-print No. quant-ph/9902068, February 1999.

54. Iliev B.Z., On metric-connection compatibility and the signature change of spacetime, → Physica Scripta, vol. 66, No. 6, pp. 401–409, 2002; IOP Publishing - IOPscience, DOI 10.1238/Physica.Regular.066a00401, URL: https://iopscience.iop.org/article/10.1238/Physica.Regular.066a00401 → http://arXiv.org e-Print archive, E-print No. gr-qc/9802058, February 1998.

55. Iliev B.Z., Fibre bundle formulation of relativistic quantum mechanics. → http://arXiv.org e-Print archive, E-print No. quant-ph/0201085, January 2002.

⇓ \underline{2003} ⟹ • 2003 · ⟶ {↙ \overleftarrow{2003} \atop ↓}

56. Iliev B.Z., On operator differentiation in the action principle in quantum field theory, → In “Trends in Complex Analysis, Differential Geometry and Mathematical Physics”, Proceedings of the 6th International Workshop on Complex Structures and Vector Fields, 3–6 September 2002, St. Knstantin resort (near Varna), Bulgaria, Editors Stancho Dimiev and Kouei Sekigava, World Scientific, New Jersey-London-Singapore-Hong Kong, 2003. pp. 76–107; → http://arXiv.org e-Print archive, E-print No. hep-th/0204003, April 2002.

57. Iliev B.Z., Normal frames for derivations and linear connections and the equivalence principle, → Journal of Geometry and Physics, vol. 45, No. 1–2, pp. 24–53, 2003. DOI No. 10.1016/S0393-0440(02)00119-5; → http://arXiv.org e-Print archive, E-print No. hep-th/0110194, October 2001.

58. Iliev B.Z., Fibre bundle formulation of relativistic quantum mechanics. I. Timedependent approach, → Physica Scripta, vol. 68, No. 1, pp. 22–31, 2003, DOI No: 10.1238/Physica.Regular.068a00022, Journal link: https://iopscience.iop.org/article/10.1238/Physica.Regular.068a00022; → http://arXiv.org e-Print archive, E-print No. quant-ph/0105056, May 2001.

⇓ \underline{2004} ⟹ • 2004 · ⟶ {↙ \overleftarrow{2004} \atop ↓}

59. Iliev B.Z., Equivalence principle in classical electrodynamics.

→ In: Has the last word been said on classical electrodynamics? – New Horizons, Editors: Andrew Chubykalo, Vladimir Onoochin, Augusto Espinoza, Roman Smirnov-Rueda, Princeton, NJ, Rinton Press, 2004, pp. 385–403, Volodymyr KrasnoholovetsISBN-13: 978-1589490369, ISBN-10: 1589490363; → http://arXiv.org e-Print archive, E-print No. gr-qc/0303002, March 2003.

60. Iliev B.Z., On momentum operator in quantum field theory, → In: Frontiers in quantum physics research, Editors Frank Columbus and Volodymyr Krasnoholovets, Nova Science Publishers, Inc., New York, 2004, \underline{List of the scientific publications of Bozhidar Z. Iliev} 8 pp. 143-156, ISBN: 1594540020, ISBN13: 9781594540028; → http://arXiv.org e-Print archive, E-print No. hep-th/0206008, June 2002.

61. Iliev B.Z., On angular momentum operator in quantum field theory, → In: Frontiers in quantum physics research, Editors Frank Columbus and Volodymyr Krasnoholovets, Nova Science Publishers, Inc., New York, 2004, pp. 129–142, ISBN: 1594540020, ISBN13: 9781594540028; → http://arXiv.org e-Print archive, E-print No. hep-th/0211153, November 2002.

62. Iliev B.Z., On conserved operator quantities in quantum field theory, → In: Frontiers in quantum physics research, Editors Frank Columbus and Volodymyr Krasnoholovets, Nova Science Publishers, Inc., New York, 2004, pp. 157–166, ISBN: 1594540020, ISBN13: 9781594540028; → http://arXiv.org e-Print archive, E-print No. hep-th/0301134, January 2003

63. Iliev B.Z., Pictures and equations of motion in quantum field theory, → In: Studies in Mathematical Physics Research, Editor Charles V. Benton, Nova Science Publishers, Inc., New York, 2004. pp. 83–125, ISBN-10: 1594540276, ISBN-13: 978-1594540271; → http://arXiv.org e-Print archive, E-print No. hep-th/0302002, February 2003

64. Iliev B.Z., Lagrangian quantum field theory in momentum picture. II. Free spinor fields → http://arXiv.org e-Print archive, E-print No. hep-th/0405008, May 1, 2004.

⇓ \underline{2005} ⟹ • 2005 · ⟶ {↙ \overleftarrow{2005} \atop ↓}

65. Iliev B.Z., On fibre bundle formulation of classical and statistical mechanics. → In “Contemporary aspects of complex analysis, differential geometry and mathematical physics” Proceedings of the 7th International Workshop on “Complex Structures and Vector Fields”, Plovdiv, Bulgaria, 31 August - 4 September 2004. Editors Stancho Dimiev and Kouei Sekigava, World Scientific, New Jersey-London-Singapore-Hong Kong, 2005. pp. 100–108; → http://arXiv.org e-Print archive, E-print No. physics/0109033, September 2001.

66. Iliev B.Z., Normal frames and linear transports along paths in vector bundles, → JP Journal of Geometry and Topology, vol. 5, No. 3, pp. 187–250, 2005; → http://arXiv.org e-Print archive, E-print No. gr-qc/9809084, September 1998.

67. Iliev B.Z., Links between connections, parallel transports, and transports along paths in differentiable fibre bundles. → International Journal of Geometric Methods in Modern Physics, vol. 2, No. 5, pp. 823–838, 2005; → Talk at the International Workshop on “Advanced Geometric Methods in Physics”, Florence, Italy, 14 – 18 April, 2005. DOI No: 10.1142/S0219887805000806; → http://arXiv.org e-Print archive, E-print No. math.DG/0504010, April 68. Iliev B.Z., Lagrangian quantum field theory in momentum picture. I. Free scalar

fields → In: Quantum Field Theory: New Researcn, Editor O. Kovras, Nova Science Publishers, Inc., New York, 2005, pp. 1–66; → http://arXiv.org e-Print archive, E-print No. hep-th/0402006, February, 2004.

69. Iliev B.Z., Lagrangian quantum field theory in momentum picture. III. Free vector fields. → http://arXiv.org e-Print archive, E-print No. hep-th/0505007, May 1, 2005.

⇓ \underline{2006} ⟹ • 2006 · ⟶ {↙ \overleftarrow{2006} \atop ↓}

70. Iliev B.Z., Handbook of Normal Frames and Coordinates.

→ Series Progress in Mathematical Physics, Vol. 42, Birkhäuser, Basel-Boston-Berlin, 2006, 468 pages, Hardcover, ISBN (ISBN-10): 3-7643-7618-X, ISBN-13: 978-3-7643-7618-5, EAN: 9783764376185. → http://arXiv.org e-Print archive, E-print No. math.DG/0610037, October 1, 2006.

71. Iliev B.Z., On the axiomatization of “parallel transport” and its generalization. → Tensor, N. S., vol. 62, No. 2, pp. 170–181, 2006; → Report presented at the 8^{th} International Conference of Tensor Society on “Differential geometry, functional and complex analysis, informatics and their applications”, Varna, Bulgaria, 22 – 26 August, 2005 → http://arXiv.org e-Print archive, E-print No. math-ph/0510005, October 1, 2005.

72. Iliev B.Z., Connection theory on differentiable fibre bundles: A concise introduction. → International Journal Mathematics and Mathematical Sciences, vol. 2006, Article No. 56397, pp. 1–51, 2006; DOI No. 10.1155/IJMMS/2006/56397; → http://arXiv.org e-Print archive, E-print No. math-ph/0510004, October 1, 2005.

73. Iliev B.Z., Normal frames for general connctions on differentiable fibre bundles → Journal of Geometry and Physics, vol. 56, No. 5, pp. 780–812, 2006, DOI No. 10.1016/j.geomphys.2005.04.018; → http://arXiv.org e-Print archive, E-print No. math.DG/0405004, May 1, 2004.

74. Iliev B.Z. and M. Rahula, Linear vector fields and exponential law.

→ http://arXiv.org e-Print archive, E-print No. math.DG/0604005, April 1, 2006.

⇓ \underline{2007} ⟹ • 2007 · ⟶ {↙ \overleftarrow{2007} \atop ↓}

75. Iliev B.Z., Lagrangian quantum field theory in momentum picture. IV. Commutation relations for free fields. → http://arXiv.org e-Print archive, E-print No. arXiv:0704.066 [hep-th], April 1, 2007.

76. Iliev B.Z., Normal frames and linear transports along paths in line bundles. Applications to classical electrodynamics. → In “Topics in contemporary differential geometry, complex analysis and mathematical physics”, Proceedings of the 8th International Workshop on Complex Structures and Vector Fields, Institute of Mathematics and Informatics, Sofia, Bulgaria, 21 – 26 August 2006, Editors Stancho Dimiev and Kouei Sekigava, World Scientific, New Jersey-London-Singapore-Hong Kong, 2007. pp. 119–129; → Report presented at the 8thv International Workshop on “Complex Structures and Vector fields”, Sofia, Bulgaria, 21–26 August 2006; → http://arXiv.org e-Print archive, E-print No. math-ph/0603002 (password: 6grgj), March 1, 2006.

⇓ \underline{2008} ⟹ • 2008 · ⟶ {↙ \overleftarrow{2008} \atop ↓}

77. Iliev B.Z., Langrangian Quantum Field theory in Momentum Picture. Free fields. → Nova Science Publishers, New York, 2008, Hardcover, 306 pages., ISBN-13: 978-1-60456-170-8, ISBN-10: 1-60456-170-X.

78. Iliev B.Z., Momentum picture of motion in Lagrangian quantum field theory, → International Journal of Theoretical Physics, Group Theory, and Nonlinear Optics, vol. 12, No. 2, pp. 103–121, 2008; → http://arXiv.org e-Print archive, E-print No. hep-th/0311003, November 2003.

79. Iliev B.Z., Affine and fundamental vector fields.

→ International journal of mathematics, game theory and algebra, vol. 17, No. 3, pp. 105–124, 2008; → http://arXiv.org e-Print archive, E-print No. math.DG/0602006, February 1, 2006.

80. Iliev B.Z., Fibre bundle formulation of relativistic quantum mechanics. II. Covariant approach, → International Journal of Theoretical Physics, Group Theory, and Nonlinear Optics, vol. 12, No. 2, pp. 85–102, 2008; → http://arXiv.org e-Print archive, E-print No. quant-ph/0107002, July 2001.

81. Iliev B.Z., Connections, parallel transports and transports along paths. → International Journal of Geometric Methods in Modern Physics, vol. 5, No. 7, pp. 1137–1162, 2008. DOI No: 10.1142/S0219887808003260; → http://arXiv.org e-Print archive, E-print No. math.DG/0504010v4, May 1, 2008.

82. Iliev B.Z., The “Lorenz gauge” is named in honour of Ludowig Valentin Lorenz! → http://arXiv.org e-Print archive, E-print No. arXiv:08.03.0047 [physics.histph], May 1, 2008.

⇓ \underline{2009} ⟹ • 2009 · ⟶ {↙ \overleftarrow{2009} \atop ↓}

83. Iliev B.Z., Heisenberg relations in the general case.

→ In ”Trends differential geometry, complex analysis and mathematical physics”, Proceedings of the 9th International Workshop on ”Complex Structures, integrability and Vector Fields”, Sofia, Bulgaria, 25-29 August 2008. Editors Kouei Sekigava, Vladimir S. Gerdjikov and Stancho Dimiev, World Scientific, New Jersey - London - Singapore - Beijing - Shanghai - Hong Kong - Taipei - Chennai, 2009, pp. 120–130 → Report presented at the 9th International Workshop on “Complex Structures, Integrability and Vector fields”, Sofia, Bulgaria, 25–29 August 2008; → http://arXiv.org e-Print archive, E-print No. arXiv:0809.0174 [math-ph], September 1, 2008.

⇓ \underline{2011} ⟹ • 2011 · ⟶ {↙ \overleftarrow{2011} \atop ↓}

84. Iliev B.Z., A mathematical base for fibre bundle formulation of Lagrangian quantum field theory. → In: International Workshop on Complex Structures, Integrability and Vector Fields, Sofia, Bulgaria, 13-17 September 2010, AIP Conference Proceedings Vol. 1340, Editors: Kouei Sekigava, Vladimir S. Gerdjikov, Yasuo Matsusita and Ivailo Mladenov, American Institute of Physics, Meilville, New York, 2011, pp. 42–52; → Report presented at the 10^{th} International Workshop on “Complex Structures, Integrability and Vector fields”, Sofia, Bulgaria, 13–17 September 2010; → http://arXiv.org e-Print archive, E-print No. arXiv:1010.3892v1 [mathph]

⇓ \underline{2012} ⟹ • 2012 · ⟶ {↙ \overleftarrow{2012} \atop ↓}

85. Iliev B.Z., Transformation laws of the components of classical and quantum fields and Heisenberg relations. → International Journal of Modern Physics A, vol. 27, No. 26, pp. 1250152-1–1250152-24, 2012. DOI No: 10.1142/S0217751X12501527; → http://arXiv.org e-Print archive, E-print No. arXiv:1203.6254 [math-ph], March 28, 2012.

⇓ \underline{2013} ⟹ • 2013 · ⟶ {↙ \overleftarrow{2013} \atop ↓}

86. Iliev B.Z., Measuring the evaluation and impact of scientific works and their authors. → http://arXiv.org e-Print archive, E-print No. arXiv:1311.6948 [cs.DL], November 27, 2013.

⇓ \underline{2014} ⟹ • 2014 · ⟶ {↙ \overleftarrow{2014} \atop ↓}

87. Iliev B.Z., Modern bibliometric indicators and achievements of authors. Modern bibliometric indicators and achievements of authors. Journal of Geometry and Symmetry in Physics, vol. 33, pp. 113–128, 2014;

⇓ \underline{2016} ⟹ • 2016 · ⟶ {↙ \overleftarrow{2016} \atop ↓}

88. Iliev B.Z., Measuring the evaluation and impact of scientic works and their authorss. Journal of Geometry and Symmetry in Physics, vol. 40, pp. 69–119, 2015 (Published March 2016);

Produced on January 5, 2025.